OpenStudy (anonymous):

Solve the quadratic equation y2 - 5y = -4 by factoring.

OpenStudy (pfenn1):

\[y^2-5y+4=0\]\[(y+a)(y+b)=0\]where\[a+b=-5\]\[ab=4\]

OpenStudy (mertsj):

\[y^2-5y+4=0 (y-4)(y-1)=0\]